# Single attosecond XUV pulse source via light-wave controlled relativistic laser-plasma interaction: Thomson Back Scattering Scheme

### v1

@@ VI Summary and Conclusion
Through comprehensive particle-in-cell simulations, we demonstrate that a carefully synthesized two-color drive pulse, combining the fundamental frequency with its second harmonic, enables the controlled generation of a single, dense RES. This RES acts as a relativistic mirror, coherently reflecting and compressing a counter-propagating source pulse into a bright, attosecond burst. The temporal shaping of the drive field is shown to be critical, governing both the formation and acceleration dynamics of the RES. Moreover, we show that by adjusting the temporal delay between the RES and the source pulse, the carrier frequency of the scattered AP can be precisely tuned. The RES longitudinal velocity can be further boosted by using standard techniques with a second reflector [39, 56] enabling generation of attosecond pulses deep into the water-window spectral range.
This tunability, combined with the demonstrated robustness of the scheme across multiple simulation geometries, underscores its experimental viability. These findings mark a significant step forward in the field of relativistic optics, highlighting the power of laser waveform synthesis in controlling ultrafast plasma dynamics. The ability to generate bright, isolated attosecond pulses with high spectral purity and tailored properties promises new avenues for attosecond spectroscopy, ultrafast solid-state physics, and high-field science. The results are also important for high density electron bunch acceleration in the few MeV regime, which have important radiobiological applications. These developments also underscore the importance of synergy between advanced target fabrication and laser waveform engineering, for enabling a new regime of precision relativistic optics at the nanoscale.
@@ Acknowledgements
ELI-ALPS is supported by the European Union and co-financed by the European Regional Development Fund (ERDF) (GINOP-2.3.6-15-2015-00001). S.K. and M.S. acknowledge the High Performance Computation (HPC) facility at ELI-ALPS.

### v2

@@ VI Summary and Conclusion
Through comprehensive particle-in-cell simulations, we demonstrate that a carefully synthesized two-color drive pulse, combining the fundamental frequency with its second harmonic, enables the controlled generation of a single, dense RES. This RES acts as a relativistic mirror, coherently reflecting and compressing a counter-propagating source pulse into a bright, attosecond burst. The temporal shaping of the drive field is shown to be critical, governing both the formation and acceleration dynamics of the RES. Moreover, we show that by adjusting the temporal delay between the RES and the source pulse, the carrier frequency of the scattered AP can be precisely tuned.
This tunability, combined with the demonstrated robustness of the scheme across multiple simulation geometries, underscores its experimental viability. These findings mark a significant step forward in the field of relativistic optics, highlighting the power of laser waveform synthesis in controlling ultrafast plasma dynamics. The ability to generate bright, isolated attosecond pulses with high spectral purity and tailored properties promises new avenues for attosecond spectroscopy, ultrafast solid-state physics, and high-field science. The results are also important for high density electron bunch acceleration in the few MeV regime, which have important radiobiological applications [54]. These developments also underscore the importance of synergy between advanced target fabrication and laser waveform engineering, for enabling a new regime of precision relativistic optics at the nanoscale.
The proposed scheme is grounded in two key recent advancements. First, the laser technology required to experimentally realize this approach is rapidly advancing [55, 56, 57], with high-intensity, waveform-controlled pulses becoming increasingly available at facilities such as ELI [58, 59, 3, 60] and other laboratories worldwide [61, 1]. Second, innovation in ultrathin target fabrication has enabled relativistic laser-plasma interactions using few-nanometer-thick large-area carbon-based targets [62] or replenishable microfluidic systems [63, 9, 64]. These targets offer a unique combination of ultrathin geometry, high repetition rate compatibility, and surface smoothness, making them ideal for precision-controlled interactions with intense laser fields. When integrated with state-of-the-art, waveform-controlled laser systems, these ultrathin targets offer a unique testbed for investigating light-wave-driven electron dynamics—a central focus of this study. The ability to tailor the temporal and spatial structure of the laser field at sub-cycle resolution opens new avenues for coherent control of plasma surfaces, efficient RES formation, and attosecond-scale electron bunch generation.
@@ Acknowledgements
The ELI ALPS project (GINOP-2.3.6-15-2015-00001) was supported by the European Union and co-financed by the European Regional Development Fund. S.K. and M.S. acknowledge the High Performance Computation (HPC) facility at ELI-ALPS.